OpenStudy (anonymous):

Find the domain of the given function. f(x)=squareroot(x+9)/(x+3)(x-7) I think it's this: x ≥ -9, x ≠ -3, x ≠ 7 But I'm not sure. i know for a fact that x definitely can't equal to 7. My other choices are: x ≠ -9, x ≠ -3, x ≠ 7 All real numbers x > 0

hero (hero):

Your answer choice is correct. We know it is not all real numbers. We also now that x is allowed to equal -9 because \(\sqrt{-9 + 9} = \sqrt{0}\) and we know we can take the square root of zero. Finally x > 0 doesn't include the entire domain and it includes x = 7 which we know is not part of the domain.

OpenStudy (anonymous):

Ohh okay! Thank you so much! It looked the most correct to me but I wanted to verify. Thanks again!